Problem 12. 4 pts Taylor Hill wants to invest $90,000 in a two-asset portfolio with a beta of 1.0. The stock XX has a beta of 2.0 while the stock YY has a beta of 0.75. How much will she invest () in each stock to maintain a portfolio beta of 1.0? (Show your calculations)
Answer #1

Let investment in XX=$x

Hence investment in YY=(90,000-x)

Portfolio beta=Respective beta*Respective investment weight

1=(x/90,000*2)+(90,000-x)/90,000*0.75

(1*90,000)=2x+67500-0.75x

90,000=2x+67500-0.75x

x=(90,000-67500)/(2-0.75)

=$18000=investment in XX

Hence investment in YY=(90,000-18000)=$72000.